# Moduli stacks of Higgs bundles on stable curves

### v2

@@ 4 The logarithmic Dolbeault moduli stack and its log symplectic structure
$(-\cap\eta_{A}):R{p_{A}}_{*}\mathcal{E}\longrightarrow(R{p_{A}}_{*}(\mathcal{E}^{\vee}))^{\vee}[-2]=(R{p_{A}}_{*}(\mathcal{E}^{\vee})[2])^{\vee}$ (4.12)
is a quasi-isomorphism of $A$-dg-modules, where $\eta_{A}:Rp_{A*}\mathcal{O}_{\mathcal{X}_{Dol,A}}\longrightarrow A[-2]$ is the derived pullback of $\eta:Rp_{*}\mathcal{O}_{\mathcal{X}_{Dol/\mathcal{S}}}\longrightarrow\mathcal{O}_{\mathcal{S}}[-2]$ under the map ${\rm Spec}\,(A)\longrightarrow\mathcal{S}$ and $(R{p_{A}}_{*}(\mathcal{E}^{\vee}))^{\vee}$ is the derived $A$-dual of $R{p_{A}}_{*}(\mathcal{E}^{\vee})$. Here
@@
corresponding to $R{q_{A}}_{*}\mathcal{E}_{A}$ (here, $E_{A}$ is sitting in degree 00). The right hand side is isomorphic to $(R{r_{A}}_{*}(\mathsf{Dol}^{\mathsf{D}}_{(E_{A},\phi_{A})}))^{\vee}$, where
$\mathsf{Dol}^{\mathbf{D}}_{(E_{A},\phi_{A})}:=[E^{\vee}_{A}\xrightarrow{-\phi^{\vee}_{A}}E^{\vee}_{A}\otimes\omega_{\mathcal{X}_{A},A}]$ (4.13)
$-2\hskip 99.58464pt-1$ (4.14)
where $\phi^{\vee}_{A}$ is the image of $\phi_{A}$ under the natural isomorphism $\mathcal{E}nd(E_{A})\cong\mathcal{E}nd(E^{\vee}_{A})$. Now since the Grothendieck-Serre dual of the complex $\mathsf{Dol}_{(E_{A},\phi_{A})}$ is the same as $\mathsf{Dol}^{\mathbf{D}}_{(E_{A},\phi_{A})}$, the morphism (4.12) is quasi-isomorphism. (We have used the notation $\mathsf{Dol}^{\mathbf{D}}_{(E_{A},\phi_{A})}$ in (4.13) to disambiguate between the ordinary dual and the Serre dual of the complex $\mathsf{Dol}_{(E_{A},\phi_{A})}$.)
Therefore, we have shown that $\mathcal{X}_{Dol/\mathcal{S}}$ is $\mathcal{O}$-compact and $\mathcal{O}$-oriented over ${\mathcal{S}}$ and from [32, Theorem 2.5], it follows that $\mathsf{Map}_{\mathcal{S}}(\mathcal{X}_{Dol/\mathcal{S}},BGL_{n}(\mathcal{O}_{\mathcal{S}}))$ has a 00-shifted relative symplectic structure over ${\mathcal{S}}$. ∎
Let us fix an integer $n$, which will represent rank in the moduli problem. Let $S$ be a spectrum of a discrete valuation ring over $\mathbb{k}$. We start with the set-up as in 3.1.1, i.e., with a semistable curve $\mathcal{X}\to S$ whose generic fiber is smooth and whose closed fiber has a single node. Let us denote the relative dualising sheaf by $\omega_{\mathcal{X}/S}$. Let $\mathfrak{M}$ denote the stack of expanded degenerations of $\mathcal{X}/S$ bounded by the integer $n$. We denote the universal curve over $\mathfrak{M}$ by $\mathfrak{X}$ (see subsection 3.2). We drop the subscript $\mathfrak{M}$ from $\mathfrak{X}_{\mathfrak{M}}$, for simplicity. We denote the relative logarithmic Dolbeault stack of $\mathfrak{X}/\mathfrak{M}$ by $\mathfrak{X}_{Dol/\mathfrak{M}}$.
@@
${\mathfrak{X}_{Dol/\mathfrak{M}}}$${\mathfrak{X}}$${\mathfrak{M}}$$\scriptstyle{q}$$\scriptstyle{p}$$\scriptstyle{r}$ (4.15)
@@
$\mathbb{T}_{M/\mathfrak{M}}\cong Rp_{*}\mathcal{E}nd\leavevmode\nobreak\ \mathcal{E}[1]\cong Rr_{*}\mathcal{C}(E,\phi)[1]$ (4.16)
@@
$\left[\mathcal{E}nd(E)\xrightarrow{[-,\phi]}\mathcal{E}nd(E)\otimes\mathsf{pr}_{\mathfrak{X}}^{*}\omega_{\mathfrak{X}/\mathfrak{M}}\right]$ (4.17)
with $\mathcal{E}nd(E)$ sitting in the degree 00. Here $\mathcal{E}$ denotes the universal sheaf on $\mathfrak{X}_{Dol/\mathfrak{M}}\times_{\mathfrak{M}}M$ and $(E,\phi)$ denotes the corresponding universal relative Higgs complex on $\mathfrak{X}\times_{\mathfrak{M}}M$. From [40, Theorem 0.3], it follows that $\mathbb{L}_{M/\mathfrak{M}}$ is a perfect complex because the map $r:\mathfrak{X}\longrightarrow\mathfrak{M}$ is a proper representable local complete intersection morphism. Also, using Grothendieck-Serre duality we can see that the complex $Rr_{*}\mathcal{C}(E,\phi)[1]$ has amplitude in $[-1,1]$. Therefore, we conclude that the morphism
@@
The derived Artin stack $\mathsf{Map}_{{}_{\mathfrak{M}}}(\mathfrak{X}_{Dol/\mathfrak{M}},BGL_{n}(\mathcal{O}_{\mathfrak{M}}))$ has a natural relative 00-shifted log-symplectic structure over $S$.
@@
$\mathbb{L}_{M/\mathfrak{L}^{0}_{S,\mathsf{lf}}}\cong\mathbb{L}_{M/\mathfrak{M}}\cong Rr_{*}\mathcal{C}(E,\phi)[1].$ (4.18)
So, the relative log-cotangent complex of the map $M\longrightarrow S$ is isomorphic to the relative cotangent complex of the map $M\longrightarrow\mathfrak{M}$. From 4.7, it follows that the stack $M$ has a 00-shifted symplectic structure relative to the stack $\mathfrak{M}$, which is, by definition 3.19, a 00-shifted log-symplectic structure on $M$ relative to $S$.
@@
$\mathbb{T}_{BGL_{n}}\wedge\mathbb{T}_{BGL_{n}}=\mathfrak{g}l_{n}[1]\wedge\mathfrak{g}l_{n}[1]\xrightarrow{(\mathsf{x},\mathsf{y})\mapsto Trace(\mathsf{x}\cdot\mathsf{y})}\mathbb{k}[2].$ (4.19)
@@
$f^{*}_{A}\mathbb{T}_{BGL_{n}}\otimes f^{*}_{A}\mathbb{T}_{BGL_{n}}=\mathcal{E}nd\leavevmode\nobreak\ \mathcal{E}_{A}[1]\otimes\mathcal{E}nd\leavevmode\nobreak\ \mathcal{E}_{A}[1]\xrightarrow{Tr:=Trace}\mathcal{O}_{\mathfrak{X}_{Dol,A}}[2]$ (4.20)
@@
${{q_{A}}_{*}(\mathcal{E}nd\leavevmode\nobreak\ E_{A}[1]\otimes\mathcal{E}nd\leavevmode\nobreak\ E_{A}[1])}$${(\mathcal{O}_{\mathcal{X}_{A}}\oplus\omega_{\mathfrak{X}_{A},A}[-1])[2]}$ (4.21)
The complex ${q_{A}}_{*}(\mathcal{E}nd\leavevmode\nobreak\ \mathcal{E}_{A}[1])=[\mathcal{E}nd\leavevmode\nobreak\ E_{A}\xrightarrow{[-,\phi_{A}]}\mathcal{E}nd\leavevmode\nobreak\ E_{A}\otimes\omega_{\mathfrak{X}_{A},A}]$ can be used to describe the pullback of the pairing in (4.21) by the natural map
${q_{A}}_{*}(\mathcal{E}nd\leavevmode\nobreak\ \mathcal{E}_{A}[1])\otimes{q_{A}}_{*}(\mathcal{E}nd\leavevmode\nobreak\ \mathcal{E}_{A}[1])\to{q_{A}}_{*}(\mathcal{E}nd\leavevmode\nobreak\ \mathcal{E}_{A}[1]\otimes\mathcal{E}nd\leavevmode\nobreak\ \mathcal{E}_{A}[1])$
@@
where the map $d_{1}:=([-,\phi],[-,\phi])$ and $d_{2}:=[-,\phi]\otimes\mathbb{1}+\mathbb{1}\otimes[-,\phi]$. Finally, by pushing forward (4.21) by the map $r_{A}$ and then composing with the orientation $\eta_{A}$ (see (4.12)), we get our log-symplectic pairing. From the above description it is evident that this log-symplectic form coincides with the log-symplectic form discussed in [9]. ∎
5 Completeness of the Hitchin map
In this section, we define the Hitchin map on the ordinary Artin stack of Gieseker-Higgs bundles $M^{cl}_{Gie}$ (see 8.1.2). We prove that the Hitchin map is complete. This result was proved in [1] for the Hitchin map as defined on the moduli scheme of stable Higgs bundles in the case where rank and degree are co-prime. We prove it here for the moduli stack, and the argument does not require us to assume that the rank and degree are co-prime.
@@ 5 Completeness of the Hitchin map
[38, p7] Working over $\operatorname{Spec}\mathbb{k}$, a morphism of Artin stacks $f:\mathcal{Y}_{1}\longrightarrow\mathcal{Y}_{2}$ is called complete if given any discrete valuation ring $A$ over $\mathbb{k}$ with fraction field $K$ and any commutative diagram
${\operatorname{Spec}\leavevmode\nobreak\ K}$${\mathcal{Y}_{1}}$${\operatorname{Spec}\leavevmode\nobreak\ A}$${\mathcal{Y}_{2}}$$\scriptstyle{f}$ (5.1)
there exists finite field extension $K\longrightarrow K^{\prime}$ with $A^{\prime}$ the integral closure of $A$ in $K^{\prime}$ and a lift over (the dotted arrow) $\operatorname{Spec}\leavevmode\nobreak\ A^{\prime}\longrightarrow\mathcal{X}$ making all the triangles in the diagram commute.
${\operatorname{Spec}\leavevmode\nobreak\ K^{\prime}}$${\operatorname{Spec}\leavevmode\nobreak\ K}$${\mathcal{Y}_{1}}$${\operatorname{Spec}\leavevmode\nobreak\ A^{\prime}}$${\operatorname{Spec}\leavevmode\nobreak\ A}$${\mathcal{Y}_{2}}$$\scriptstyle{f}$ (5.2)
@@
We recall the following setup from before. Let $S$ be a spectrum of a discrete valuation ring over $\mathbb{k}$. We start with the set-up as in 3.1.1, i.e., with a semistable curve $\mathcal{X}\to S$ whose generic fiber is smooth and whose closed fiber has a single node. Let us denote the relative dualising sheaf by $\omega_{\mathcal{X}/S}$. Let $\mathfrak{M}$ denote the stack of expanded degenerations of $\mathcal{X}/S$ bounded by the integer $n$. We denote the universal curve over $\mathfrak{M}$ by $\mathfrak{X}$ (see subsection 3.2). We drop the subscript $\mathfrak{M}$ from $\mathfrak{X}_{\mathfrak{M}}$, for simplicity. We denote the relative logarithmic Dolbeault stack of $\mathfrak{X}/\mathfrak{M}$ by $\mathfrak{X}_{Dol/\mathfrak{M}}$.
@@
Now consider the substack $M^{cl}_{Gie}\subseteq\tau_{0}(M)\subseteq M$ in the classical truncation $\tau_{0}(M)$ of $M$ parametrizing Higgs bundles whose underlying vector bundle satisfies the Gieseker conditions. Namely, it is globally generated on the fibers of the map $X_{k}\to X_{0}$ and whose pushforward to $X_{0}$ is torsion-free. See 8.1.2 for a detailed description. Note also that for semi-stable families $\mathfrak{X}/S$ of arithmetic genus greater than or equal to $2$, the morphism $\tau_{0}(M)\to S$ is l.c.i. and therefore $\tau_{0}(M)\cong M$. In particular, the derived structure on $M$ plays a role only in genus zero and one.
@@
The morphism $h|_{M^{cl}_{Gie}}:M^{cl}_{Gie}\longrightarrow B$ is complete.
@@
Let $T:=\operatorname{Spec}A$ be a spectrum of a discrete valuation ring over $\mathbb{k}$ with function field $K$. Let us denote $T^{o}:=\operatorname{Spec}K$. Suppose we are given a commutative diagram
@@
We need to prove that this diagram factorizes through a map $T\to M^{cl}_{Gie}$. Let $TFH(\mathcal{X}/S)$ denote the ordinary Artin stack of families of torsion-free $\omega_{\mathcal{X}/S}$-valued Higgs pairs on the family of curves $\mathcal{X}/S$ (see Definition 8.1.1). Then there is a morphism
$\theta:M^{cl}_{Gie}\rightarrow TFH(\mathcal{X}/S)$
$(\pi:\mathcal{X}^{mod}\to\mathcal{X},\mathcal{E},\phi)\to(\pi_{*}\mathcal{E},\pi_{*}\phi)$
given by the pushforward of Higgs bundles under some modification $\mathcal{X}^{\textsf{mod}}\rightarrow\mathcal{X}$ of Gieseker curves, as defined in 3.2. Therefore, from (5.3) we get the following diagram
@@
Using [15, Proposition 6] and [24, Lemma 6.5], we have a map $T\to TFH(\mathcal{X}/S)$ making the triangle commute:

### v3

@@ 4. The logarithmic Dolbeault moduli stack and its log symplectic structure
(4.12) $(-\cap\eta_{A}):R{p_{A}}_{*}\mathcal{E}\longrightarrow(R{p_{A}}_{*}(\mathcal{E}^{\vee}))^{\vee}[-2]=(R{p_{A}}_{*}(\mathcal{E}^{\vee})[2])^{\vee}$
is a quasi-isomorphism of $A$-dg-modules, where $\eta_{A}:Rp_{A*}\mathcal{O}_{\mathcal{X}_{Dol,A}}\longrightarrow A[-2]$ is the derived pullback of $\eta:Rp_{*}\mathcal{O}_{\mathcal{X}_{Dol/\mathcal{S}}}\longrightarrow\mathcal{O}_{\mathcal{S}}[-2]$ under the map ${\rm Spec}\,(A)\longrightarrow\mathcal{S}$ and $(R{p_{A}}_{*}(\mathcal{E}^{\vee}))^{\vee}$ is the derived $A$-dual of $R{p_{A}}_{*}(\mathcal{E}^{\vee})$. Here,
@@
corresponding to $R{q_{A}}_{*}\mathcal{E}_{A}$ (here, $E_{A}$ is in degree 0). The right hand side is isomorphic to $)(R{r_{A}}_{*}(\mathsf{Dol}^{\mathsf{D}}_{(E_{A},\phi_{A})}$$)^{\vee}$, where
(4.13) $\mathsf{Dol}^{\mathbf{D}}_{(E_{A},\phi_{A})}:=[E^{\vee}_{A}\xrightarrow{-\phi^{\vee}_{A}}E^{\vee}_{A}\otimes\omega_{\mathcal{X}_{A},A}]$
(4.14) $-2\hskip 99.58464pt-1$
where $\phi^{\vee}_{A}$ is the image of $\phi_{A}$ under the natural isomorphism $\mathcal{E}\mathsf{nd}(E_{A})\cong\mathcal{E}\mathsf{nd}(E^{\vee}_{A})$. Now since the Grothendieck-Serre dual of the complex $\mathsf{Dol}_{(E_{A},\phi_{A})}$ is the same as $\mathsf{Dol}^{\mathbf{D}}_{(E_{A},\phi_{A})}$, the morphism (4.12) is quasi-isomorphism. (We have used the notation $\mathsf{Dol}^{\mathbf{D}}_{(E_{A},\phi_{A})}$ in (4.13) to disambiguate between the ordinary dual and the Serre dual of the complex $\mathsf{Dol}_{(E_{A},\phi_{A})}$.)
Therefore, we have shown that $\mathcal{X}_{Dol/\mathcal{S}}$ is $\mathcal{O}$-compact and $\mathcal{O}$-oriented over $\mathcal{S}$} and from [29, Theorem 2.5], it follows that $\mathsf{Map}_{\mathcal{S}}(\mathcal{X}_{Dol/\mathcal{S}},BGL_{n}(\mathcal{O}_{\mathcal{S}}))$ has a 0-shifted relative symplectic structure over $\mathcal{S}$}. ∎
Remark 4.8.
The above theorem is more broadly applicable to any given family of reduced Gorenstein projective algebraic curves $(\mathcal{X}/S$. The main focus of the theorem is on two properties: $\mathcal{O}$-compactness and $\mathcal{O}$-orientedness of the family $\mathcal{X}_{\text{Dol}}/S$. These properties hold in this context due to the correspondences discussed in Remark 4.6.
Firstly, these correspondences apply to any geometric Derived stacks, as noted in [32, Section 5.1]. In particular, they will apply to $\mathcal{X}_{\text{Dol}}/S$ for any given family of projective algebraic curves $\mathcal{X}/S$. However, the proof also depends on the fact that the dualizing sheaf is locally free and that Grothendieck-Serre duality holds. For this reason, the Gorenstein condition is essential.
Fix an integer $n$, which will represent rank in the moduli problem. Let $S$ be the spectrum of a discrete valuation ring on $\mathbb{k}$. We start with the set-up as in §3.1.1, i.e., with a semistable curve $\mathcal{X}\to S$ whose generic fiber is smooth and whose closed fiber has a single node. Let us denote the relative dualizing sheaf by $\omega_{\mathcal{X}/S}$. Let $\mathfrak{M}$ denote the stack of expanded degenerations of $\mathcal{X}/S$ bounded by the integer $n$. We denote the universal curve over $\mathfrak{M}$ by $\mathfrak{X}$ (see subsection 3.2). We drop the subscript $\mathfrak{M}$ from $\mathfrak{X}_{\mathfrak{M}}$, for simplicity. We denote the relative logarithmic Dolbeault stack of $\mathfrak{X}/\mathfrak{M}$ by $\mathfrak{X}_{Dol/\mathfrak{M}}$.
@@
(4.15) $\mathfrak{X}_{Dol/\mathfrak{M}}$$\mathfrak{X}$$\mathfrak{M}$$\scriptstyle{q}$$\scriptstyle{p}$$\scriptstyle{r}$
@@
(4.16) $\mathbb{T}_{M/\mathfrak{M}}\cong Rp_{*}\mathcal{E}\mathsf{nd}\penalty 10000\ \mathcal{E}[1]\cong Rr_{*}\mathcal{C}(E,\phi)[1]$
@@
(4.17) $\left[\mathcal{E}\mathsf{nd}(E)\xrightarrow{[-,\phi]}\mathcal{E}\mathsf{nd}(E)\otimes\mathsf{pr}_{\mathfrak{X}}^{*}\omega_{\mathfrak{X}/\mathfrak{M}}\right]$
with $\mathcal{E}\mathsf{nd}(E)$ sitting in the degree 0. Here $\mathcal{E}$ denotes the universal sheaf on $\mathfrak{X}_{Dol/\mathfrak{M}}\times_{\mathfrak{M}}M$ and $(E,\phi)$ denotes the corresponding universal relative Higgs complex on $\mathfrak{X}\times_{\mathfrak{M}}M$. From [42, Theorem 0.3], it follows that $\mathbb{L}_{M/\mathfrak{M}}$ is a perfect complex because the map $r:\mathfrak{X}\longrightarrow\mathfrak{M}$ is a proper representable local complete intersection morphism. Also, using Grothendieck-Serre duality we can see that the complex $Rr_{*}\mathcal{C}(E,\phi)[1]$ has amplitude in [1,1][-1,1]. Therefore, we conclude that the morphism
@@
Remark 4.10.
It is important to note that for the proof of the proposition to be valid, we required two key criteria regarding the family of curves $\mathfrak{X}\to\mathfrak{M}$: properness and the local complete intersection property (e.g. nodal or cuspidal singularities). These criteria ensure that the derived pushforward of any perfect complex remains perfect and that the relative version of Grothendieck-Serre duality holds for such a family of curves.
@@
The derived Artin stack $\mathsf{Map}_{{}_{\mathfrak{M}}}(\mathfrak{X}_{Dol/\mathfrak{M}},BGL_{n}(\mathcal{O}_{\mathfrak{M}}))$ has a natural relative 0-shifted log-symplectic structure over $S$. Moreover, this form coincides with the relative log-symplectic form described in [10].
@@
(4.18) $\mathbb{L}_{M/\mathfrak{L}^{0}_{S,\mathsf{lf}}}\cong\mathbb{L}_{M/\mathfrak{M}}\cong Rr_{*}\mathcal{C}(E,\phi)[1].$
So, the relative log-cotangent complex of the map $M\longrightarrow S$ is isomorphic to the relative cotangent complex of the map $M\longrightarrow\mathfrak{M}$. From Theorem 4.7, it follows that the stack $M$ has a 0-shifted symplectic structure relative to the stack $\mathfrak{M}$, which is, by definition 3.23, a 0-shifted log-symplectic structure on $M$ relative to $S$.
@@
(4.19) $\mathbb{T}_{BGL_{n}}\wedge\mathbb{T}_{BGL_{n}}=\mathfrak{g}l_{n}[1]\wedge\mathfrak{g}l_{n}[1]\xrightarrow{(\mathsf{x},\mathsf{y})\mapsto Trace(\mathsf{x}\cdot\mathsf{y})}\mathbb{k}[2].$
@@
(4.20) $f^{*}_{A}\mathbb{T}_{BGL_{n}}\otimes f^{*}_{A}\mathbb{T}_{BGL_{n}}=\mathcal{E}\mathsf{nd}\penalty 10000\ \mathcal{E}_{A}[1]\otimes\mathcal{E}\mathsf{nd}\penalty 10000\ \mathcal{E}_{A}[1]\xrightarrow{Tr:=Trace}\mathcal{O}_{\mathfrak{X}_{Dol,A}}[2]$
@@
(4.21) ${q_{A}}_{*}(\mathcal{E}\mathsf{nd}\penalty 10000\ E_{A}[1]\otimes\mathcal{E}\mathsf{nd}\penalty 10000\ E_{A}[1])$$(\mathcal{O}_{\mathcal{X}_{A}}\oplus\omega_{\mathfrak{X}_{A},A}[-1])[2]$}
The complex ${q_{A}}_{*}(\mathcal{E}\mathsf{nd}\penalty 10000\ \mathcal{E}_{A}[1])=[\mathcal{E}\mathsf{nd}\penalty 10000\ E_{A}\xrightarrow{[-,\phi_{A}]}\mathcal{E}\mathsf{nd}\penalty 10000\ E_{A}\otimes\omega_{\mathfrak{X}_{A},A}]$ can be used to describe the pullback of the pairing in (4.21) by the natural map
${q_{A}}_{*}(\mathcal{E}\mathsf{nd}\penalty 10000\ \mathcal{E}_{A}[1])\otimes{q_{A}}_{*}(\mathcal{E}\mathsf{nd}\penalty 10000\ \mathcal{E}_{A}[1])\to{q_{A}}_{*}(\mathcal{E}\mathsf{nd}\penalty 10000\ \mathcal{E}_{A}[1]\otimes\mathcal{E}\mathsf{nd}\penalty 10000\ \mathcal{E}_{A}[1])$
@@
where the map $d_{1}:=([-,\phi_{A}],[-,\phi_{A}])$ and $d_{2}:=[-,\phi_{A}]\otimes\mathbb{1}+\mathbb{1}\otimes[-,\phi_{A}]$. Finally, by pushing forward (4.21) by the map $r_{A}$ and then composing with the orientation $\eta_{A}$ (see (4.12)), we get our log-symplectic pairing. From the above description it is evident that this log-symplectic form coincides with the log-symplectic form discussed in [10]. ∎
Remark 4.12.
Note that the proof of the first statement of the above Theorem depends on the following crietria: (i) there is a log-structures on the space $\mathfrak{M}$, (ii) the map $M\to\mathfrak{M}$ is quasismooth, (iii) there is a relative 0-shifted symplectic form on $M\to\mathfrak{M}$ and finally the fact that (iv) $\mathbb{L}_{\mathfrak{M}/\mathfrak{L}^{0}_{S,\mathsf{lf}}}\cong 0$. With these four hypothesis, we have shown that there is a relative 0-shifted shifted symplectic form on $M\to S$. The second statement of the theorem is about the specific situtation appearing from a one parameter degeneration of a family of stable curves.
5. Completeness of the Hitchin map
In this section, we define the Hitchin map on the ordinary Artin stack of Gieseker-Higgs bundles $M^{cl}_{Gie}$ (see A.1). We prove that the Hitchin map is complete. This result was proved in [2] for the Hitchin map as defined on the moduli scheme of stable Higgs bundles in the case where rank and degree are co-prime. We prove it here for the moduli stack, and the argument does not require us to assume that the rank and degree are co-prime.
@@ 5. Completeness of the Hitchin map
[40, p7] Working over $\operatorname{Spec}\mathbb{k}$, a morphism of Artin stacks $f:\mathcal{Y}_{1}\longrightarrow\mathcal{Y}_{2}$ is called complete if given any discrete valuation ring $A$ over $\mathbb{k}$ with fraction field $K$ and any commutative diagram
(5.1) $\operatorname{Spec}\penalty 10000\ K$$\mathcal{Y}_{1}$$\operatorname{Spec}\penalty 10000\ A$$\mathcal{Y}_{2}$$\scriptstyle{f}$
there exists finite field extension $K\longrightarrow K^{\prime}$ with $A^{\prime}$ the integral closure of $A$ in $K^{\prime}$ and a lift over (the dotted arrow) $\operatorname{Spec}\penalty 10000\ A^{\prime}\longrightarrow\mathcal{X}$ making all the triangles in the diagram commute.
(5.2) $\operatorname{Spec}\penalty 10000\ K^{\prime}$$\operatorname{Spec}\penalty 10000\ K$$\mathcal{Y}_{1}$$\operatorname{Spec}\penalty 10000\ A^{\prime}$$\operatorname{Spec}\penalty 10000\ A$$\mathcal{Y}_{2}$$\scriptstyle{f}$
@@
Remark 5.2.
Notice that from the definition it is obvious, in particular, that the fibers of a complete morphism of Artin stacks are complete because the morphism in the diagram 5.1 is allowed to factor through the map $\operatorname{Spec}\penalty 10000\ A\to\operatorname{Spec}k\to\mathcal{Y}_{2}$. The notion of completeness is the same as the notion of uiversal closedness of the morphism of stacks ([41, Proposition 26.20.6 (Valuative criterion of universal closedness).]).
We recall the following setup from before. Let $S$ be a spectrum of a discrete valuation ring over $\mathbb{k}$. We start with the set-up as in §3.1.1, i.e., with a semistable curve $\mathcal{X}\to S$ whose generic fiber is smooth and whose closed fiber has a single node. Let us denote the relative dualising sheaf by $\omega_{\mathcal{X}/S}$. Let $\mathfrak{M}$ denote the stack of expanded degenerations of $\mathcal{X}/S$ bounded by the integer $n$. We denote the universal curve over $\mathfrak{M}$ by $\mathfrak{X}$ (see subsection 3.2). We drop the subscript $\mathfrak{M}$ from $\mathfrak{X}_{\mathfrak{M}}$, for simplicity. We denote the relative logarithmic Dolbeault stack of $\mathfrak{X}/\mathfrak{M}$ by $\mathfrak{X}_{Dol/\mathfrak{M}}$.
@@
Now consider the substack $M^{cl}_{Gie}\subseteq\tau_{0}(M)\subseteq M$ in the classical truncation $\tau_{0}(M)$ of $M$ parameterizing Higgs bundles whose underlying vector bundle satisfies the Gieseker conditions. That is, it is generated globally on the fibers of the map $X_{k}\to X_{0}$ and its push-forward to $X_{0}$ is torsion-free. See A.1 for a detailed description.
@@
The morphism $h|_{M^{cl}_{Gie}}\colon M^{cl}_{Gie}\longrightarrow B$ is complete.
@@
Let $T=\operatorname{Spec}A$ be the spectrum of a discrete valuation ring over $\mathbb{k}$, with fraction field $K$. Denote the generic point by $T^{o}=\operatorname{Spec}K$. Suppose that we are given a commutative diagram
@@
We must show that this diagram extends to a morphism $T\to M^{cl}_{Gie}$. Let $\mathsf{TFH}(\mathcal{X}/S)$ be the ordinary Artin stack of families of torsion-free $\omega_{\mathcal{X}/S}$-valued Higgs pairs on $\mathcal{X}/S$ (see Definition A.1.2). There is a natural morphism
$\theta\colon M^{cl}_{Gie}\longrightarrow\mathsf{TFH}(\mathcal{X}/S),\qquad(\pi\colon\mathfrak{X}\to\mathcal{X},\ \mathcal{E},\ \phi)\mapsto((\pi_{*}\mathcal{E}),\ (\pi_{*}\phi)).$
This is induced by pushforward along the modification $\mathfrak{X}\to\mathcal{X}$. Applying $\theta$ to the original diagram gives a commutative triangle
(5.4) $T^{o}$$\mathsf{TFH}(\mathcal{X}/S)$}T{T}
By [20, Proposition 6] and [26, Lemma 6.5], there exists a morphism $T\to\mathsf{TFH}(\mathcal{X}/S)$ making the triangle commute. Now consider the surface $\mathcal{X}_{T}:=\mathcal{X}\times_{S}T$. There are two cases:
Case 1. The map $T\to S$ is not faithfully flat. Then $\mathcal{X}_{T}\cong X_{0}\times T$. The normalisation of $\mathcal{X}_{T}$ is $\widetilde{X}_{0}\times T$. Let $f\colon\widetilde{X}_{0}\to X_{0}$ be the normalisation map. By [9, Proposition 3.2], the pullback of the generic Higgs bundle extends to a good generalized parabolic Hitchin pair on $\widetilde{X}_{0}\times T$, which in turn gives a torsion-free Hitchin pair on $\mathcal{X}_{T}$.
Case 2. The map $T\to S$ is faithfully flat. Then $\mathcal{X}_{T}$ is a normal surface with an isolated singularity of type $\mathbb{k}[[x,y,t]]/(xy-t^{d})$. Let $i\colon\mathcal{X}_{T^{o}}\hookrightarrow\mathcal{X}_{T}$ be the inclusion of the generic fibre. Suppose we have a Higgs bundle $(\mathcal{E}^{o},\phi^{o})$ on $\mathcal{X}_{T^{o}}$. Any vector bundle on the generic fibre extends to a torsion-free sheaf $\mathcal{F}$ on $\mathcal{X}_{T}$. We wish to extend the Higgs field to $\mathcal{F}\to\mathcal{F}\otimes\omega_{\mathcal{X}_{T}/T}$. Consider the diagram
@@
By [26, Lemma 6.5], the Higgs field extends over $\mathcal{X}_{T}$ except possibly at the node. The composite map $\mathcal{F}\to(i_{*}\mathcal{E}^{o}/\mathcal{F})\otimes\omega_{\mathcal{X}_{T}/T}$ vanishes generically on the closed fibre. Thus the obstruction lies in the torsion part of $i_{*}\mathcal{E}^{o}/\mathcal{F}$ (viewed as a sheaf on its support $X_{0}$). Restricting the top sequence to the closed fibre gives
(5.6) $0\longrightarrow\mathsf{Tor}^{1}_{{}_{\mathcal{X}_{T}}}(i_{*}\mathcal{E}^{o}/{\mathcal{F}},\mathcal{O}_{X_{0}})\longrightarrow\mathcal{F}|_{X_{0}}\longrightarrow(i_{*}\mathcal{E}^{o})|_{X_{0}}\longrightarrow i_{*}\mathcal{E}^{o}/{\mathcal{F}}\longrightarrow 0.$
Since $\mathsf{Tor}^{1}_{{}_{\mathcal{X}_{T}}}(i_{*}\mathcal{E}^{o}/{\mathcal{F}},\mathcal{O}_{X_{0}})=(i_{*}\mathcal{E}^{o}/\mathcal{F})\otimes\mathcal{O}_{\mathcal{X}_{T}}(-X_{0})$ and $\mathcal{O}_{\mathcal{X}_{T}}(-X_{0})$ is locally free, the sheaf $i_{*}\mathcal{E}^{o}/\mathcal{F}$ is pure (torsion-free on its support) if and only if the Tor term is pure. But the Tor term is a subsheaf of $\mathcal{F}|_{X_{0}}$, hence torsion-free on $X_{0}$. This implies the obstruction vanishes, and the Higgs field extends everywhere.
In both cases we obtain a commutative diagram